Two physics students are separated 6.80 m and are creating a standing wave interference pattern along a slinky consisting of 5 loops. If the students are vibrating their ends of the slinky with a frequency of 3.10 Hz, calculate the speed of the wave energy.


1
Expert's answer
2021-01-29T04:29:25-0500
"L=\dfrac{1}{2}n\lambda=\dfrac{5}{2}\lambda=2.5\lambda."

From this formula we can find the wavelength of the wave travelling in the slinky:


"\lambda=\dfrac{L}{2.5}=\dfrac{6.8\ m}{2.5}=2.72\ m."

Finally, we can find the speed of the wave from the wave speed formula:


"v=f\lambda=3.1\ Hz\cdot2.72\ m=8.43\ \dfrac{m}{s}."